# Planetarium[Uncaptioned image]: A Rigorous Benchmark for Translating Text to Structured Planning Languages

### v1

@@ 4.2 The Dataset
The Planetarium dataset includes 132,037 text-to-PDDL pairs across 13 tasks, derived from the Blocks World and Gripper domains Seipp et al. (2022); McDermott (2000).
@@
We focus on these two domains as they are commonly used in other works Valmeekam et al. (2023a); Liu et al. (2023), but are nevertheless challenging. Each text-to-PDDL pair consists of a prompt detailing the initial and goal states, along with the corresponding ground truth PDDL. (See Appendix D for examples.) Table 1 shows a breakdown of the Planetarium data by text description abstractness and number of propositions. Each instance is a ground truth PDDL problem description and text description pair.
Table 1: A breakdown of the Planetarium dataset by the level of abstractness of the text description and number of propositions in the ground truth PDDL problem (size). Each instance is a ground truth PDDL problem description and text description pair.
Domain Total Abstractness Num. of Ground Truth Propositions (Size)
Explicit to Abstract to
Explicit Abstract Explicit Abstract 1-20 21-40 41-60 61-80 >80
Blocks World 92,350 23,033 23,087 23,086 23,144 1,012 10,765 50,793 26,316 3,464
Gripper 39,687 10,632 9,518 10,313 9,224 379 2,112 9,412 25,346 2,438
We vary the data along two dimensions: abstractness (explicit vs abstract) and size. Explicit planning problem text descriptions correspond directly to propositions found in the problem PDDL (e.g., “block 1 is on block 2”). Abstract text descriptions instead summarize a state (e.g., “all blocks are in a single tower”). Since our text descriptions describe both initial and goal states, each can be either an abstract or explicit description, leading to four possible categories: explicit to explicit, explicit to abstract, abstract to explicit, and abstract to abstract. We measure the size of a problem by the number of propositions listed in the ground truth problem PDDL. Larger problems typically pose greater challenges for LLMs.
To obtain interesting problems for each domain, we crafted a set of tasks instead of randomly generating instances. Below are descriptions of each domain and their corresponding tasks.
Blocks World. This domain involves using a robotic hand to manipulate a set of blocks arranged in various configurations on a table. We have designed the following tasks for this domain:
  • Stack: Arrange all the blocks into a single tower.
  • Unstack: Dismantle any tower so that all blocks are separately placed on the table.
  • Holding a Block: Similar to unstacking, but one block remains held by the robot.
  • Staircase: Create stacks of increasing sizes from one upwards until no blocks are left.
  • Equal Towers: Arrange the blocks into towers of equal height.
  • Swap: With two towers, swap the base blocks and leave the rest of the structure unchanged.
  • Invert: Invert each stack so the entire sequence of blocks is reversed.
  • Towers: Build towers to match a specified list of heights.
Gripper. This domain features a robot equipped with grippers, tasked with transporting balls from one room to another. We have designed the following tasks for this domain:
  • Single Room: Move all balls to one room.
  • Evenly Distribute: Position the balls so each room has the same number of balls.
  • Arbitrarily Distribute: Move the balls according to a predefined list that specifies the number of balls per room.
  • Swap Rooms: Exchange the balls between two rooms.
  • Move to Max: Move all balls to the room currently containing the most balls.
  • Move to Min: Move all balls to the room with the least number of balls.
@@ 5 Evaluating LLMs on Planetarium
Figure 2: Performance of various models on the Planetarium test set.
As an initial snapshot of the field’s current state, we evaluate several API-access and open-weight language models on Planetarium in both zero-shot and fine-tuned settings. We find that while powerful models like GPT-4o can often generate valid PDDL problems (in the sense that some plan solves them), they are rarely correct. This result underscores the need for Planetarium’s rigorous approach to evaluation of PDDL generation. The code to recreate the entire evaluation is available at https://github.com/BatsResearch/planetarium.
Models. We evaluate one API-access model (GPT-4o) and three open-weight models (Mistral v0.3 7B Instruct Jiang et al. (2023), and Gemma 1.1 IT 2B & 7B Team et al. (2024)).
@@
We evaluate models on the Planetarium test set, which we built by randomly splitting the Planetarium dataset with an $80/20$ train/test split. Models are prompted with the natural language description of the task along with the respective domain PDDL. We measure three metrics: the number of syntactically parseable problems generated, the number of solve-able problems, and the number of semantically correct problems. We say a model output is parseable if a PDDL parser supporting :strips can parse a valid PDDL problem from a substring in the output and if it can be converted into our graph representation. A problem is solve-able if it is parseable and there exists a plan that can be applied to the initial scene that results in the goal scene. Due to the size and complexity of some of the problems in our dataset, a generalized classical planner cannot always reliably and quickly return solutions. Therefore, we write specialized planners that work for all problems in our dataset, and generally, all validly defined blocksworld and gripper domain problems, except a few invalid edge cases (e.g., one block on top of two blocks at a time, holding two blocks, etc.). To evaluate the correctness of model outputs, they must be first parseable and solve-able. Then, we use our PDDL equivalence algorithm to verify equivalence to our ground truth PDDL. We found our equivalence algorithm to take on average $12$ms per example to compute on an M2 Apple Silicon laptop with batch parallelization.
Fine-Tuning. We fine-tuned the open-weight models using QLoRA with a rank of 16, adhering to the hyperparameter recommendations for small models provided by the original authors Dettmers et al. (2023). Models were loaded in 4-bit precision and were trained over the training set for a single epoch. Fine-tuning for each model used either two NVIDIA GeForce RTX 3090 GPUs or two NVIDIA A6000 GPUs, operating with data parallelization for approximately 15 hours. We truncate the longest $5\%$ of our training dataset due to GPU memory constraints. Experiments were conducted using a GPU cluster at the Center for Computation and Visualization, Brown University. Additional experiment details are provided in Appendix F.
@@ Results.
The performance of GPT-4o, Mistral v0.3 7B Instruct, and Gemma 1.1 IT 2B & 7B on the Planetarium test set is shown in Figure 2. We report the percentage of generated plans that were parseable and correct and report results from zero-shot and fine-tuned settings.
Figure 3: Breakdown of GPT-4o zero-shot performance at varying levels of abstraction.
Analysis and Discussion. All models exhibited poor zero-shot performance, with GPT-4o achieving the highest accuracy. A breakdown of GPT-4o’s zero-shot performance, the only model to achieve significant zero-shot performance, shows abstract task descriptions are harder to translate than explicit ones, and fully explicit task descriptions make generating parseable PDDL code easier (Figure 3).
Fine-tuning improved performance across all open-weight models, with Mistral v0.3 7B Instruct achieving the highest accuracy. More experiments and analyses are shown in Appendix G.
6 Conclusion, Limitations, and Future Work
Planetarium is a new benchmark for assessing the ability of LLMs to translate natural language descriptions of planning problems into PDDL. A potential societal impact of this research is ensuring the correctness of translating natural language into structured planning languages. If this translation method becomes widespread but its evaluation remains inaccurate, systems could produce misleading or misaligned results that could cause harm if acted upon. Planetarium highlights the importance of assessing the correctness of translations. We achieve this by reasoning about PDDL semantics and the inherent structure of classical planning problems.
Planetarium has a few important limitations: Planetarium currently only supports the Blocks World and Gripper domains. While these domains have been popular for studying LLMs and their relation to planning, incorporating more expressive domains in the future will widen the scope of Planetarium. This benchmark is also currently restricted to the STRIPS subset of PDDL. Extending it to support more expressive subsets of PDDL will allow us to evaluate more complex, real-world planning problems such as non-deterministic, temporal, and numeric domains. Our work reveals generating semantically correct structured planning language descriptions to be a challenging task for language models, with models like GPT-4o achieving $35.1\%$ zero-shot accuracy but producing valid, seemingly correct descriptions $82.2\%$ of the time. We hope that Planetarium will drive progress on hybrid approaches combining LLMs and classic planners, setting a standard for the evaluation of such tasks.
Acknowledgments and Disclosure of Funding
@@ Appendix A Proof of Theorem 1
isEquivalent($P_{a}$, $P_{b}$, False) returns True if and only if the PDDL problem files $P_{a}$ and $P_{b}$ represent equivalent planning problems under Definition 2.
@@
To prove this proposition, we need to prove each direction of the biconditional. First, we will prove the forward implication. There are two possible ways the algorithm can return True: Either in the isIsomorphic check (line 18) or in fastEquivalence (line 6).
  1. Case 1:
    When isIsomorphic($\text{p}_{a}$, $\text{p}_{b}$) returns True, we know that there exists a bijection $\varphi$ from vertices $V_{p_{a}}$ to vertices $V_{p_{b}}$. Since $\varphi$ is a type-preserving bijection, it means the two graphs share a connectivity structure where all types and attributes match.
@@ 1. Case 1:
    To prove Property 1, we can build the set $L^{a}$ for each planning problem by taking each predicate in the PDDL domain and instantiating it with all possible combinations of objects from the problem PDDL $P_{a}$. Then, we can construct the set of all possible states of the problem by taking the power set $S^{a}=2^{L^{a}}$. For each proposition $p\in s$ in each state $s\in S^{a}$, we apply the bijection $f$ to show that we get $S^{b}$.
    $S$ =$=\{\{f(p):p\in s\}:s\in S^{a}\}$ (1)
    $S$ =$=\{\{\langle\varphi(q^{a}),\varphi(o_{1}^{a}),\varphi(o_{2}^{a}),\ldots,\varphi(o_{k}^{a})\rangle:p\in s\}:s\in S^{a}\}$ (2)
    $S$ =$=\{\{\langle q^{b},o_{1}^{b},o_{2}^{b},\ldots o_{k}^{b}\rangle:p\in s\}:s\in S^{a}\}$ (3)
    $S$ =$=\{\{p_{b}:p\in s\}:s\in S^{a}\}$ (4)
    $S$ $=S^{b}$ (5)
    To prove Property 2, we go through each proposition $p$ in the initial state $s_{i}^{a}$ of the problem PDDL $P_{a}$, and apply the bijection $f$ to show that we get $s_{i}^{b}$.
    $s_{i}$ =$=\{f(p_{a}):p_{a}\in s_{i}^{a}\}$ (6)
    $s_{i}$ =$=\{\langle\varphi(q^{a}),\varphi(o_{1}^{a}),\varphi(o_{2}^{a}),\ldots,\varphi(o_{k}^{a})\rangle:p_{a}\in s_{i}^{a}\}$ (7)
    $s_{i}$ =$=\{\langle q^{b},o_{1}^{b},o_{2}^{b},\ldots o_{k}^{b}\rangle:p_{a}\in s_{i}^{a}\}$ (8)
    $s_{i}$ =$=\{p_{b}:p_{a}\in s_{i}^{a}\}$ (9)
    $s_{i}$ $=s_{i}^{b}$ (10)
    To prove Property 3, we first observe that the fullySpecify function adds all possible trivial propositions to the goal scene graph. By definition, trivial propositions are those that are known not to change the underlying planning problem when added or removed. We can use the same bijection $f$ constructed from $\varphi$. We don’t need to add or remove non-trivial predicates, as doing so would change the planning problem by definition. This implies that the augmented goal state represents any other possible goal state in $S_{g}$. With this, we can use the bijection $f$ built from $\varphi$ to apply it to any state in $S_{g^{a}}$ and show that we get $S_{g^{b}}$.
    $S_{g}$ =$=\{\{f(p):p\in s\}:s\in S_{g^{a}}\}$ (11)
    $S_{g}$ =$=\{\{\langle\varphi(q^{a}),\varphi(o_{1}^{a}),\varphi(o_{2}^{a}),\ldots,\varphi(o_{k}^{a})\rangle:p\in s\}:s\in S_{g^{a}}\}$ (12)
    $S_{g}$ =$=\{\langle q^{b},o_{1}^{b},o_{2}^{b},\ldots o_{k}^{b}\rangle:p\in s\}:s\in S_{g^{a}}\}$ (13)
    $S_{g}$ =$=\{\{p_{b}:p\in s\}:s\in S_{g^{a}}\}$ (14)
    $S_{g}$ $=S_{g^{b}}$ (15)
  2. Case 2:
@@
    In Case 1 we have proven that when isIsomorphic($p_{a}$, $p_{b}$) returns True, the planning problems $P_{a}$ and $P_{b}$ are equivalent under the bijection $f$. Since $p_{a}=\texttt{join(}s_{i,a},g_{a}^{\star}\texttt{)}$ and $p_{b}=\texttt{join(}s_{i,b},g_{b}^{\star}\texttt{)}$, we can let $g_{a}^{\star}=g_{a}$ and $g_{b}^{\star}=g_{b}$. This is because, by definition, trivial propositions do not change the underlying planning problem when added or removed, and in this case, we are only removing trivial propositions. Hence, by the proof of Case Case 1: we conclude that this case has also to be true.
Now, we will prove the backward implication. In this case, we will assume that there exists a bijection $f$ that follows properties 12, and 3 of Definition 2. By Lemma 1, we know that $f$ comes from $\varphi$, and since both are bijections, we can recover $\varphi$ by doing the inverse process. We need to prove that these conditions lead to isIsomorphic (line 18) or fastEquivalence (line 6) returning True.
Since there exists a bijection $\varphi$ that comes from $f$, the number of objects must be the same for planning problems $P_{a}$ and $P_{b}$, as stated in Property 1. With Property 2, we know that this bijection $\varphi$ must also make the initial scenes isomorphic. Assume that the sets of goal propositions have the same size. By using Property 3, we know that the bijection $\varphi$ would make fastEquivalence (line 6) return True.
Conversely, assume that the sets of goal propositions have different sizes. In this case, canFastEquivalence would not allow fastEquivalence to execute. In the case of isIsomorphic (line 18), we know that $p_{a}=\texttt{join(}s_{i,a}g_{a}^{\star}\texttt{)}$ and $p_{b}=\texttt{join(}s_{i,b},g_{b}^{\star}\texttt{)}$. Since we know that we can construct the bijection $\varphi$ that maps vertices from the bijection $f$, we also know that since $\varphi$ comes from $f$, this bijection respects initial states (Property 2) and goal states (Property 3). Therefore, isIsomorphic($p_{a}$, $p_{b}$) must return True.
@@
isEquivalent($P_{a}$, $P_{b}$, True) returns True if and only if $P_{a}$ represents a planning problem that is equivalent (under Definition 2) to some planning problem represented by $P_{b}$ after a permutation of the objects in its goal state.
@@
To prove this proposition, we need to prove each direction of the biconditional. First, we will prove the forward implication. There are two possible ways the algorithm can return True: Either in fastEquivalence (line 6) or in isIsomorphic($s_{i,a}$, $s_{i,b}$) $\wedge$ isIsomorphic($g^{\star}_{a}$, $g^{\star}_{b}$).
To prove the case where fastEquivalence (line 6) returns True, we can use the same argument as in Case Case 1: of the forward implication in the proof of Proposition 1. This is because the placeholder being True does not affect the fastEquivalence function.
@@
Since the underlying mappings of $f$ and $h$ can be different, we need a way to combine them into a single bijection. One way to construct this bijection is to permute the mappings of $f$ until they fit the mappings of $h$. This implies that there are many pairs of $f$ and $h$, and each pair can be combined into a global bijection using function composition $f\circ h$. For each of these permutations, we can then use the arguments in Case Case 2: of the proof of Proposition 1 to show that these pairs hold all three properties of Definition 2.
Now, we will prove the backward implication. First, since fastEquivalence remains the same, we do not need to prove anything new about it; we can simply reuse the previous results.

### v2

@@ 4.2 The Dataset
The Planetarium dataset includes 145,918 text-to-PDDL pairs derived from the Blocks World, Gripper, and Floor Tile domains  Seipp et al. (2022); McDermott (2000). Our dataset captures 25 unique initial and goal state configurations describing 73 different tasks. In the Blocks World domain, a robotic hand manipulates blocks arranged in various configurations on a table, with 5 predicates and 4 actions. The Gripper domain features a robot with grippers transporting balls between rooms, with 7 predicates and 3 actions. While versions of Blocks World and Gripper domains exist with :typing, we choose to use the IPC standard versions, which do not. We simplified the standard Floor Tile domain Seipp et al. (2022), where robots are tasked with painting tiles, by removing the left and down predicates and relying on the right and up predicates to check relative positions. We further relax the domain by removing the clear predicate (which introduces path constraints) and the free-color predicate (which IPC problems did not use and is not modified or required by any actions in the original domain). Finally, we added paint-left and paint-right actions, which reduced edge cases for the domain, resulting in a domain with 6 predicates and 9 actions (compared to 10 predicates and 7 actions in the IPC version). More details about the dataset can be found in Appendix D.
Table 2: Examples of abstract and concrete descriptions for each domain in the benchmark.
Domain Abstract Explicit
blocksworld
You have 3 blocks, stacked into 2 towers of heights
1, 2, and your arm is empty. Your goal is to invert
each individual stack of blocks, such that the block
that in each tower that was originally on the bottom
will be on the top.
Your arm is empty. b1 is clear. b1 is on the table.
b2 is clear. b2 is on b3. b3 is on the table. Your goal
is to have the following: Your arm should be empty.
b3 should be clear. b3 should be on b2. b2 should be
on the table. b1 should be clear. b1 should be on the
table.
gripper
You have 2 rooms, 2 balls, and 2 grippers. 1 balls are
distributed across the same number of grippers, and
the rest are in the first room. The robby is in the first
room. Your goal is to gather all balls into one room.
You have 2 rooms, 2 balls, and 2 grippers. 1 balls
are distributed across the same number of grippers,
and the rest are in the first room. The robby is in the
first room. Your goal is to have the following: gripper1
should be free. gripper2 should be free. ball1 should be
at room1. ball2 should be at room1.
floor-tile
You have 1 robots, 2 colors, and 2 unpainted tiles
arranged in a grid with 1 rows and 2 columns. The
first robot is at the top-left corner, and has the first
color. All colors are available. Your goal is to paint
all the tiles with the same color.
You have 1 robot. You have 2 tiles. You have 2 colors.
Tile tile2 is to the right of tile tile1. The robot robot1
has color color1. The robot robot1 is at tile tile1. Color color1
is available. Color color2 is available. Your goal is to have
the following: Tile tile1 should be painted with color color1.
Tile tile2 should be painted with color color1.
@@
We focus on these three domains commonly used in other works Valmeekam et al. (2023a); Liu et al. (2023) but are nevertheless challenging. Each entry in our dataset is a text-to-PDDL pair consisting of a text description detailing the initial and goal states and the corresponding ground truth PDDL. (See Appendix H for examples.) The composition of the Planetarium dataset is presented in Table 1, which categorizes the data according to text description abstractness and the number of propositions. The table further delineates the train/test split, illustrating the distribution of data reserved for evaluation purposes. For a comprehensive overview of the task descriptions and the specific allocation of tasks between training and test sets, readers are directed to Appendix D. The dataset is procedurally generated: we handcraft templates for each task configuration, which we combine with one another to generate problems at scale.
We vary the data along two dimensions: abstractness (explicit vs. abstract) and size. Explicit planning problem text descriptions correspond directly to propositions found in the problem PDDL (e.g., “block 1 is on block 2”). Abstract text descriptions instead summarize a state (e.g., “all blocks are in a single tower”).
Since our text descriptions contain both initial and goal states, each can be either an abstract or explicit description. This leads to four possible abstractness categories: explicit to explicit, explicit to abstract, abstract to explicit, and abstract to abstract.
We measure the size of a problem by the number of propositions listed in the ground truth problem PDDL. Larger problems typically pose greater challenges for LLMs.
Figure 3: Performance of various models on the Planetarium test set.
@@ 5 Evaluating LLMs on Planetarium
As an initial snapshot of the field’s current state, we evaluate several API-access and open-weight language models on Planetarium in both zero-shot and fine-tuned settings. We find that while powerful models like GPT-4o can often generate valid PDDL problems (in the sense that some plan solves them), they are rarely correct. This result underscores the need for Planetarium’s rigorous approach to evaluation of PDDL generation. The code to recreate the entire evaluation is available at github.com/BatsResearch/planetarium.
Models.
We evaluate two API-access models, GPT-4o and o1-mini, and we evaluate three open-weight models before and after fine-tuning: Gemma 2 2B IT, Gemma 2 9B IT, and Gemma 2 27B IT Team (2024). Details regarding fine-tuning can be found in Appendix F.
@@
We evaluate models on the Planetarium test set, which consists of heldout configurations from the Blocks World and Gripper domains and the entirety of the Floor Tile domain. Models are prompted with the natural language description of the task along with the respective domain PDDL. We record three metrics for the generated problems: the number of parseable problems, the number of solvable problems, and the number of correct problems. We say a model output is parseable if a PDDL parser supporting :strips can extract a valid PDDL problem from a substring in the output and if it can be converted into our graph representation. A problem is solvable if it is parseable and a plan can be applied to the initial scene that results in the goal scene. Due to the size and complexity of some of the problems in our dataset, a generalized classical planner cannot always reliably and quickly return solutions. Instead, we built specialized planners for Blocks World and Gripper problems that work on all problems in our dataset and generally all validly defined blocksworld and gripper domain problems except for a few invalid edge cases (e.g., one block on top of two blocks at a time, holding two blocks, etc.). We use Fast Downward Helmert (2006) for floor-tile problems. We then validate all plans with VAL Howey et al. (2004). We then feed parseable and solvable problems into our PDDL equivalence algorithm to verify equivalence to our ground truth PDDL and thereby determine if it is correct. We found our equivalence algorithm to take, on average, 12ms per example to compute on an M2 Apple Silicon laptop with batch parallelization.
Figure 4: Breakdown of zero-shot performance by domain for Gemma 2 27B IT, GPT-4o, and o1-mini.
@@ Results.
The performance of Gemma 2 2/9/27B IT, GPT-4o, and o1-mini on the Planetarium test set are shown in Figure 3. We report the percentage of generated plans that were parseable, solvable, and correct from zero-shot and fine-tuned settings, averaged across all domains. For o1-mini, we randomly sample 250 problems per domain to evaluate due to prohibitive costs. In the zero-shot setting, o1-mini and GPT-4o performed the best, followed closely by Gemma 27B IT. For all models, the percentage of generated PDDL that is semantically correct significantly lags behind the percentage of generated PDDL that is parseable and solvable.
Fine-tuning improved performance across all open-weight models, with fine-tuned Gemma 2 27B IT achieving the highest accuracy of all tested models at 34.2%.
Takeaways
We find two key takeaways from the type of mistakes these language models tended to make. First, a significant portion of errors come from models “ignoring” the domain in context and referring to an incorrect domain, likely from its prior training corpus. We can see this by the large increase in accuracy when ignoring :typing, shown in Figure 4: our Blocks World and Gripper domains do not have typing, yet when generating PDDL, these models confidently and consistently add object types.
Second, there seems to be a strong correlation between a model’s performance on a certain domain and the domain’s frequency in the wild: Gripper and Blocks World are often used as introductory PDDL examples, while Floor Tile is less common. However, Gripper, and to a lesser extent, Blocks World, is often defined with :typing. This may explain the difference in performance between Blocks World and Gripper across models, as well as the difference in performance between enforcing versus not enforcing :typing.
Succinctly, we see a trend when relaxing the :typing requirement that models find problems in less common domains harder, accounting for everything else.
6 Conclusion
Planetarium is a new benchmark for assessing the ability of LLMs to translate natural language descriptions of planning problems into PDDL. Our work reveals that language models find it challenging to generate semantically correct structured planning language descriptions. Models like GPT-4o can often produce valid, seemingly correct descriptions (94.4%) when, in reality, only a small fraction (24.8%) are genuinely correct. We hope that Planetarium will drive progress on hybrid approaches combining LLMs and classic planners, setting a standard for evaluating such tasks.
Acknowledgements
@@
Limitations
Planetarium has a few important limitations: Planetarium currently only supports the Blocks World, Gripper, and Floor Tile domains. While these domains have been popular for studying LLMs and their relation to planning, incorporating more expressive domains in the future will widen the scope of Planetarium. This benchmark is also currently restricted to the STRIPS subset of PDDL. Extending it to support more expressive subsets of PDDL will allow us to evaluate more complex, real-world planning problems such as non-deterministic, temporal, and numeric domains.
Potential Impacts
A potential societal impact of this research is ensuring the correctness of translating natural language into structured planning languages. If this translation method becomes widespread but its evaluation remains inaccurate, systems could produce misleading or misaligned results that could cause harm if acted upon. Planetarium highlights the importance of assessing the correctness of translations. We achieve this by reasoning about PDDL semantics and the inherent structure of classical planning problems.
@@ Appendix A Proof of Theorem 1
Equivalent($P_{a}$, $P_{b}$, False) returns True if and only if the PDDL problem files $P_{a}$ and $P_{b}$ represent equivalent planning problems under Definition 2.
@@
To prove this proposition, we need to prove each direction of the biconditional. First, we will prove the forward implication. There are two possible ways the algorithm can return True: Either in the isIsomorphic check (line 16) or in fastEquivalent (line 5).
Case 1: When isIsomorphic($\text{p}_{a}$, $\text{p}_{b}$) returns True, we know that there exists a bijection $\varphi$ from vertices $V_{p_{a}}$ to vertices $V_{p_{b}}$. Since $\varphi$ is a type-preserving bijection, it means the two graphs share a connectivity structure where all types and attributes match.
@@
To prove Property 1, we can build the set $L^{a}$ for each planning problem by taking each predicate in the PDDL domain and instantiating it with all possible combinations of objects from the problem PDDL $P_{a}$. Then, we can construct the set of all possible states of the problem by taking the power set $S^{a}=2^{L^{a}}$. For each proposition $p\in s$ in each state $s\in S^{a}$, we apply the bijection $f$ to show that we get $S^{b}$:
$S=\{\{f(p):p\in s\}:s\in S^{a}\}$
=$=\{\{\langle\varphi(q^{a}),\varphi(o_{1}^{a}),\ldots,\varphi(o_{k}^{a})\rangle:p\in s\}:s\in S^{a}\}$
=$=\{\{\langle q^{b},o_{1}^{b},\ldots,o_{k}^{b}\rangle:p\in s\}:s\in S^{a}\}$
=$=\{\{p_{b}:p\in s\}:s\in S^{a}\}$
$=S^{b}.$
To prove Property 2, we go through each proposition $p$ in the initial state $s_{i}^{a}$ of the problem PDDL $P_{a}$, and apply the bijection $f$ to show that we get $s_{i}^{b}$:
$s_{i}=\{f(p_{a}):p_{a}\in s_{i}^{a}\}$
=$=\{\langle\varphi(q^{a}),\varphi(o_{1}^{a}),\ldots,\varphi(o_{k}^{a})\rangle:p_{a}\in s_{i}^{a}\}$
=$=\{\langle q^{b},o_{1}^{b},\ldots,o_{k}^{b}\rangle:p_{a}\in s_{i}^{a}\}$
=$=\{p_{b}:p_{a}\in s_{i}^{a}\}$
$=s_{i}^{b}.$
To prove Property 3, we first observe that the fullySpecify function adds all possible trivial propositions to the goal scene graph. By definition, trivial propositions are those that are known not to change the underlying planning problem when added or removed. We can use the same bijection $f$ constructed from $\varphi$. We don’t need to add or remove non-trivial predicates, as doing so would change the planning problem by definition. This implies that the augmented goal state represents any other possible goal state in $S_{g}$. With this, we can use the bijection $f$ built from $\varphi$ to apply it to any state in $S_{g^{a}}$ and show that we get $S_{g^{b}}$:
$S_{g}=\{\{f(p):p\in s\}:s\in S_{g^{a}}\}$
=$=\{\{\langle\varphi(q^{a}),\varphi(o_{1}^{a}),\ldots,\varphi(o_{k}^{a})\rangle:p\in s\}:s\in S_{g^{a}}\}$
=$=\{\langle q^{b},o_{1}^{b},\ldots o_{k}^{b}\rangle:p\in s\}:s\in S_{g^{a}}\}$
=$=\{\{p_{b}:p\in s\}:s\in S_{g^{a}}\}$
$=S_{g^{b}}.$
Finally, to prove Property 4, we observe that Equivalent($P_{a}$, $P_{b}$, False) assumes that the domain files are shared between $P_{a}$ and $P_{b}$. As a result, any bijection $\varphi$ found by the algorithm must preserve the equality of actions across both problems, since the shared domain enforces consistency in their action sets.
@@
In Case 1 we have proven that when isIsomorphic($p_{a}$, $p_{b}$) returns True, the planning problems $P_{a}$ and $P_{b}$ are equivalent under the bijection $f$. Since $p_{a}=\texttt{join(}s_{i,a},g_{a}^{\star}\texttt{)}$ and $p_{b}=\texttt{join(}s_{i,b},g_{b}^{\star}\texttt{)}$, we can let $g_{a}^{\star}=g_{a}$ and $g_{b}^{\star}=g_{b}$. This is because, by definition, trivial propositions do not change the underlying planning problem when added or removed, and in this case, we are only removing trivial propositions. Hence, by the proof of Case 1 we conclude that this case has also to be true.
Now, we will prove the backward implication. In this case, we will assume that there exists a bijection $f$ that follows properties 12, and 3 of Definition 2. By Lemma 1, we know that $f$ comes from $\varphi$, and since both are bijections, we can recover $\varphi$ by doing the inverse process. We need to prove that these conditions lead to isIsomorphic (line 16) or fastEquivalent (line 5) returning True.
Since there exists a bijection $\varphi$ that comes from $f$, the number of objects must be the same for planning problems $P_{a}$ and $P_{b}$, as stated in Property 1. With Property 2, we know that this bijection $\varphi$ must also make the initial scenes isomorphic. Assume that the sets of goal propositions have the same size. By using Property 3, we know that the bijection $\varphi$ would make fastEquivalent (line 5) return True.
Conversely, assume that the sets of goal propositions have different sizes. In this case, canDoFast would not allow fastEquivalent to execute. In the case of isIsomorphic (line 16), we know that $p_{a}=\texttt{join(}s_{i,a}g_{a}^{\star}\texttt{)}$ and $p_{b}=\texttt{join(}s_{i,b},g_{b}^{\star}\texttt{)}$. Since we know that we can construct the bijection $\varphi$ that maps vertices from the bijection $f$, we also know that since $\varphi$ comes from $f$, this bijection respects initial states (Property 2) and goal states (Property 3). Furthermore, the algorithm does not modify any actions in the domain file, and since a single domain is shared across all problems, the actions remain consistent between them (Property 4). Therefore, isIsomorphic($p_{a}$, $p_{b}$) must return True.
@@
Equivalent($P_{a}$, $P_{b}$, True) returns True if and only if $P_{a}$ represents a planning problem that is equivalent (under Definition 2) to some planning problem represented by $P_{b}$ after a permutation of the objects in its goal state.
@@
To prove this proposition, we need to prove each direction of the biconditional. First, we will prove the forward implication. There are two possible ways the algorithm can return True: Either in fastEquivalent (line 5) or in isIsomorphic($s_{i,a}$, $s_{i,b}$) $\wedge$ isIsomorphic($g^{\star}_{a}$, $g^{\star}_{b}$) (line 12).
To prove the case where fastEquivalent (line 5) returns True, we can use the same argument as in Case 1 of the forward implication in the proof of Proposition 1. This is because the placeholder being True does not affect the fastEquivalent function.
@@
Since the underlying mappings of $f$ and $h$ can be different, we need a way to combine them into a single bijection. One way to construct this bijection is to permute the mappings of $f$ until they fit the mappings of $h$. This implies that there are many pairs of $f$ and $h$, and each pair can be combined into a global bijection using function composition $f\circ h$. For each of these permutations, we can then use the arguments in Case 2 of the proof of Proposition 1 to show that these pairs hold all three properties of Definition 2.
Now, we will prove the backward implication. First, since fastEquivalent remains the same, we do not need to prove anything new about it; we can simply reuse the previous results.